# Neural Internal Model Control: Learning a Robust Control Policy via Predictive Error Feedback

### v1

@@ I Introduction
In this paper, we propose a general and flexible framework that integrates model-based and RL-based control. By simplifying the model to focus on the robot’s main body and treating its dynamics as a rigid transformation, we bypass the high-dimensional nonlinearity of complex robot systems and environmental interactions. This simplification enables us to use an explicit predictive model without requiring parameter tuning, establishing a foundation for robust closed-loop control. We compute the predictive error—defined as the discrepancy between plant and model responses—as a feedback signal to the policy, enhancing robustness and generalizability through informative feedback and extensive training. Our framework, shown in Fig. 2, can be easily integrated with other learning algorithms by incorporating this feedback loop. To demonstrate the effectiveness and versatility of our method, we apply it to two prevalent robot morphologies: quadrotors, which demand precise 3D control, and quadrupedal robots, which require adaptive locomotion to navigate complex terrain interactions. Empirical results indicate that our framework can surpass recent state-of-the-art control methods in both domains. Furthermore, we conduct a real-world deployment on a quadrotor with various rope-suspended payloads to illustrate its robustness in sim-to-real transfer.
@@ II-A Classical control and disturbance estimation
Classical adaptive controllers require estimating system parameters in a closed-loop manner to adjust control actions in real-time. $\mathcal{L}_{1}$ adaptive control has been widely used to estimate force disturbances in quadrotors [22, 23]. Neural-Fly [18] incorporated pre-trained representations to enhance rapid online adaptation through deep learning, achieving stable flight in previously unseen wind conditions. Internal Model Control (IMC) [6, 3] is a general structure for robust control that uses a predefined model to predict system responses and estimate disturbances. However, applying IMC to more complex and nonlinear systems is challenging due to the difficulty of crafting reliable models for such systems. Long et al. [13] further proposed Hybrid Internal Model, in which they leverage a learning-based internal model with unsupervised contrastive learning. The learned model serves as a predictive encoder for estimating the system responses in the latent space. Despite their advantages, these methods either rely on explicit disturbance estimation or are heavily influenced by prediction accuracy, which can significantly affect performance if compromised.
@@ II-B Learning-based adaptive control
With the rapid advancement of highly parallel simulators [17, 24, 25], reinforcement learning (RL) methods are increasingly demonstrating powerful capabilities to achieve robust control, including quadrotors [26, 20] and legged robots [11, 12, 13, 14, 15]. Unlike classical model-based control, these learning-based methods often use model-free RL algorithms like PPO [27] to train neural controllers in large-scale randomized environments. Koryakovskii et al. [28] proposed to incorporate an RL-baesd policy to generate residual actions to compensate nonlinear model predictive control in model-mismatch cases. Rapid Motor Adaptation (RMA) [12] is a pure RL-based methods, which predicts environmental parameters in the latent space using state-action histories through a teacher-student training procedure. This approach has been extended to quadrotors [26], bipedal robots [29], and humanoid robots [15]. A follow-up study found that a dual-history architecture, leveraging both short-term and long-term input/output history, achieves superior performance with a straightforward one-stage training process [14]. Conversely, DATT [20] incorporates an $\mathcal{L}_{1}$ adaptation law [21] to estimate forces applied to quadrotors directly, which are then used as policy inputs. Despite its effectiveness with precise parameter tuning, this method is sensitive to handcrafted settings and challenging to generalize to complex systems.
@@ III-A Internal Model Control
Internal Model Control (IMC) is a powerful structure for robust control, offering significant advantages over classical feedback control, particularly in managing model uncertainties and external disturbances. Unlike classical feedback control, which relies on real-time error correction based on the difference between desired and actual outputs, IMC incorporates a predictive model of the process, allowing it to anticipate and compensate for disturbances more effectively. The structures of classical feedback control and IMC are illustrated in Fig. 3, where $P$ and $\tilde{P}$ represent the plant and internal models, respectively. $Q$ is the IMC controller, and $C$ is the classical feedback controller. The reference signal, disturbance, and system output are denoted by $r$, $d$, and $y$, respectively. In IMC, the predicted response $\tilde{y}$ is subtracted from the actual response $y$, and the difference, $\tilde{d}=\tilde{y}-y$, is used as the input signal to the IMC controller. Generally, $\tilde{d}$ is non-zero due to modeling errors and unknown disturbances not accounted for in the model. Consequently, IMC can provide more robust performance by continuously adjusting the control input according to the discrepancy between the predicted and actual responses.
While the design procedure for IMC is straightforward when the plant system is easy to model [3], it becomes significantly more complex and challenging for systems with nonlinear dynamics or when accurate models are difficult to obtain [30, 31]. This complexity arises because methods like linearization or approximation often fail to capture the true behavior of nonlinear systems, leading to performance degradation. Such a dependency on precise modeling limits its practical applications in robotic systems, such as quadrotors with cable-suspended payloads and legged robots on complex terrains, where achieving a robust and accurate system model is inherently challenging.
To address these challenges, our method combines the strengths of the IMC structure with large-scale RL training. A simplified predictive model in IMC provides a rough but stable estimate of the system’s response. Instead of relying solely on precise system modeling, our approach uses large-scale randomized simulations to learn a robust control policy that compensates for inaccuracies and simplified-out complexities. Through extensive training, the RL algorithm optimizes the control policy conditioned on predictive errors, enhancing robustness and performance in real-world scenarios where exact modeling is impractical. The predictive error reveals signals about external disturbances and internal characteristics of nonlinear dynamics, enabling more informative control decisions and effective policy learning.
@@ III-B Design of a Simplified Predictive Model
The predictive model is crucial in our framework, providing the baseline needed to detect and compensate for system discrepancies and disturbances. Designing an effective predictive model involves two main challenges: accurately estimating the system response and ensuring the model’s predictions closely match actual system behavior. Accurate estimation is essential for the feedback loop to reflect the true robot state, while alignment between predictions and real responses is critical for reliable performance under varying conditions.
@@
$\tilde{\mathbf{q}}_{k+1}$ $=\mathbf{q}_{k}\otimes\mathbf{q}_{\Delta\theta},$ (5)
$\tilde{\boldsymbol{\omega}}_{k+1}$ $=\boldsymbol{\omega}_{k}+\dot{\boldsymbol{\omega}}_{k}\Delta t.$ (6)
We have established an explicit method to predict the next body state, assuming the required motion can be represented by commands at the acceleration or velocity level. Then, given the predicted body state $\tilde{\mathbf{x}}_{k+1}=\left[\tilde{\mathbf{p}}_{k+1},\tilde{\mathbf{q}}_{k+1},\tilde{\mathbf{v}}_{k+1},\tilde{\boldsymbol{\omega}}_{k+1}\right]$ and the actual body state $\mathbf{x}_{k+1}$, we compute the predictive error $\tilde{\mathbf{d}}_{k+1}$, which is then fed into the policy to provide effective feedback signals.
$\tilde{\mathbf{d}}_{k+1}=\left[\Delta\tilde{\mathbf{p}}_{k+1},1-\cos{\tilde{\Theta}},\Delta\tilde{\mathbf{v}}_{k+1},\Delta\tilde{\boldsymbol{\omega}}_{k+1}\right]$ (7)
Here, $\Delta\tilde{\mathbf{p}}_{k+1}$, $\Delta\tilde{\mathbf{v}}_{k+1}$, and $\Delta\tilde{\boldsymbol{\omega}}_{k+1}$ represent the differences in position, velocity, and angular velocity, respectively. We use the cosine distance to quantify the difference between orientations, where $\tilde{\Theta}$ is the angle difference between the two quaternions.
@@ III-C1 Quadrotors
Traditional quadrotor controllers often use multi-layered cascaded structures [38], offering flexible options for integrating neural controllers. Recent research shows that using collective thrust and bodyrate (CTBR) as the action space for neural controllers provides significant advantages [39, 20, 9]. CTBR-based quadrotor controllers exhibit strong resilience to dynamics mismatches, effective domain transfer, and high agility and performance. This action space design is inherently compatible with our framework. We can directly use the bodyrate command from the policy to update the rotation, while the linear acceleration is computed using the collective thrust to translate the robot. Consequently, the predictive error, calculated from the previous policy output and the latest successive body states, can be seamlessly integrated into the current policy input without further modifications to the learning algorithm.
@@ III-C2 Quadrupedal robots
Unlike quadrotors, legged robots, particularly quadrupedal and bipedal robots, have garnered significant attention for RL-based controllers due to their higher control dimensions and the complexities of modeling interactions with diverse terrains [8, 11, 12, 14, 19]. Typically, a neural controller for legged robots generates target actuator positions to achieve the desired body velocity, including forward velocity ($v_{x}$), lateral velocity ($v_{y}$), and yaw rate ($\omega_{z}$). To integrate our predictive model into the control loop, we use the reference velocity command as the model input instead of the high-dimensional actuator motions. By differentiating two consecutive reference velocity commands, we obtain the linear acceleration in x and y axes, and the angular acceleration for the yaw angle. In this way, we can apply the rigid transformation by setting other acceleration terms to zero, ensuring seamless integration into the existing policy structure. This method is compatible with other algorithms by simply adding a plug-and-play feedback loop to augment policy inputs.
@@ IV Experimental Results
TABLE II: Simulation results for tracking errors (unit: m) with quadrotors, averaged over five random seeds. Bold values denote the best results. Blue cells indicate performance under random wind fields, with percentages in brackets representing the relative change in error compared to static conditions. T and S refer to the teacher and student models for RMA, respectively, while DATT with GT indicates the use of ground-truth disturbance inputs instead of $\mathcal{L}_{1}$ estimates.
@@
Train MPPI+RL 0.087 0.157 (+80.46%) 0.056 0.118 (+110.71%) 0.103 0.175 (+69.90%) 0.105 0.164 (+56.19%)
@@
Eval MPPI+RL 0.224 0.321 (+43.30%) 0.171 0.251 (+46.78%) 0.262 0.364 (+38.93%) 0.265 0.358 (+35.09%)
@@ IV-B Experiments on Quadrotors
We compare our method for quadrotors against four baselines: MPPI+RL [28], PPO [27], RMA [12], DATT [20], using two variants each for RMA and DATT to enable deeper analyses. All methods use the same PPO training parameters. For MPPI+RL, MPPI serves as a sampling-based nonlinear MPC method, with the RL policy compensating the MPPI outputs. We set the sampling number to 4096 and the rollout horizon to 20 in MPPI’s sampling procedure. PPO employs a three-layer MLP policy and a temporal convolutional network (TCN) to process future waypoints in the reference trajectory, which is inherited by all other methods. In addition, RMA adopts a teacher-student training framework, with RMA(T) as the teacher model and RMA(S) as the student. RMA(T) integrates dynamic parameters and ground-truth force disturbances, while RMA(S) uses a TCN to approximate the teacher’s latent state from state-action histories. In contrast, DATT utilizes an $\mathcal{L}_{1}$ disturbance estimator and tracking error inputs to improve trajectory-tracking robustness. To separate the $\mathcal{L}_{1}$ estimator’s effect, we also include a privileged variant, DATT(GT), which directly uses ground-truth force disturbances in place of the estimator.
As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics. Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator. However, its performance declines noticeably with unexpected dynamics in evaluation. Comparing DATT(GT) and DATT($\mathcal{L}_{1}$) reveals that the learned RL policy generalizes well when provided with ground-truth inputs, suggesting that the performance drop is primarily due to the handcrafted $\mathcal{L}_{1}$ estimator, which struggles to generalize across different systems. On the other hand. the teacher-student training approach in RMA results in the student policy underperforming compared to the teacher policy and shows instability under unseen systems, highlighting its unreliability for learning-based prediction.
@@ IV-C Experiments on Quadrupedal Robots
For quadrupedal robots, we compare our methods with PPO [27] and HIM [13]. PPO serves as a standard baseline in continuous control for quadrupeds, while HIM leverages unsupervised contrastive learning to train a predictive encoder for the next latent state and body velocity. Our method enhances the PPO policy by integrating predictive error feedback. Additionally, we evaluate a variant of our method, adding predictive error feedback to both the inputs of the HIM policy and its predictive encoder during training and evaluation.
As shown in Tab. III, our method demonstrates superior performance, achieving relative improvements over HIM ranging from approximately 26% to 44% across various test conditions. The combination of our framework with HIM also shows improvements over the original HIM. Interestingly, the predictive encoder in HIM shows a slight degradation compared to standard PPO when combined with our NeuralIMC structure. We hypothesize that this is because the predictive error provides sufficient signal for robust control, rendering the additional impact of predictive encoding negligible. This suggests that predictive error feedback alone is sufficiently informative for enhancing control robustness, simplifying the neural control pipeline, and facilitating integration with other learning algorithms.

### v2

@@ I Introduction
In this paper, we propose a new learning-based control methods which leverages a predictive error feedback to enhance control robustness, inspired by model-based control. By simplifying the model to focus on the robot’s main body and treating its dynamics as a rigid transformation, we bypass the high-dimensional nonlinearity of complex robot systems and environmental interactions. This simplification enables us to use an explicit predictive model without requiring parameter tuning, establishing a foundation for robust closed-loop control. We compute the predictive error—defined as the discrepancy between plant and model responses—as a feedback signal to the policy, enhancing robustness and generalizability through informative feedback and extensive training. Our framework, shown in Fig. 2, can be easily integrated with other learning algorithms by incorporating this feedback loop. To demonstrate the effectiveness and versatility of our method, we apply it to two prevalent robot morphologies: quadrotors, which demand precise 3D control, and quadrupedal robots, which require adaptive locomotion to navigate complex terrain interactions. Empirical results indicate that our framework can surpass recent state-of-the-art control methods in both domains. Furthermore, we conduct a real-world deployment on a quadrotor with various rope-suspended payloads to illustrate its robustness in sim-to-real transfer.
@@ II-A Classical control and disturbance estimation
Classical adaptive controllers require estimating system parameters in a closed-loop manner to adjust control actions in real-time. $\mathcal{L}_{1}$ adaptive control has been widely used to estimate force disturbances in quadrotors [22, 23]. Neural-Fly [18] incorporated pre-trained representations to enhance rapid online adaptation through deep learning, achieving stable flight in previously unseen wind conditions. Internal Model Control (IMC) [6, 3] is a general structure for robust control that uses a predefined model to predict system responses and estimate disturbances. However, applying IMC to more complex and nonlinear systems is challenging due to the difficulty of crafting reliable models for such systems. Hybrid Internal Model (HIM) further leverages a learning-based internal model that employs unsupervised contrastive learning, with the learned model acting as a predictive encoder for estimating system responses in the latent space, as proposed by Long et al. [13]. Despite their advantages, these methods either rely on explicit disturbance estimation or are heavily influenced by prediction accuracy, which can significantly affect performance if compromised.
@@ II-B Learning-based adaptive control
With the rapid advancement of highly parallel simulators [17, 24, 25], reinforcement learning (RL) methods are increasingly demonstrating powerful capabilities to achieve robust control, including quadrotors [26, 20] and legged robots [11, 12, 13, 14, 15]. Unlike classical model-based control, these learning-based methods often use model-free RL algorithms like PPO [27] to train neural controllers in large-scale randomized environments. Koryakovskii et al. [28] proposed to incorporate an RL-baesd policy to generate residual actions to compensate nonlinear model predictive control in model-mismatch cases. Rapid Motor Adaptation (RMA) [12] is a pure RL-based methods, which predicts environmental parameters in the latent space using state-action histories through a teacher-student training procedure. This approach has been extended to quadrotors [26], bipedal robots [29], and humanoid robots [15]. A follow-up study found that a dual-history architecture, leveraging both short-term and long-term input/output history, achieves superior performance with a straightforward one-stage training process [14]. Conversely, DATT [20] incorporates an $\mathcal{L}_{1}$ adaptation law [21] to estimate forces applied to quadrotors directly, which are then used as policy inputs. CAL [28] integrates RL with nonlinear model predictive control (NMPC) by employing RL to generate residual actions that augment NMPC’s outputs. Although both DATT and CAL have proven effective with precise parameter tuning, they remain sensitive to handcrafted settings in their classical control components, limiting their generalizability to more complex systems.
@@ III-A Internal Model Control
Internal Model Control (IMC) is a robust control framework that outperforms classical feedback control in handling model uncertainties and disturbances. Unlike feedback control, which corrects errors in real time, IMC uses a predictive model to anticipate and compensate for disturbances. Fig. 3 illustrates their structures, where $P$ and $\tilde{P}$ are the plant and internal models, $Q$ is the IMC controller, and $C$ is the classical feedback controller. The reference, disturbance, and output are denoted by $r$, $d$, and $y$, respectively. IMC computes the difference $\tilde{d}=\tilde{y}-y$ between predicted and actual responses, adjusting control inputs accordingly for enhanced robustness.
IMC is straightforward when the plant is easily modeled [3] but becomes challenging for nonlinear systems where approximations fail to capture true dynamics [30, 31]. This reliance on precise models limits its use in robotics, such as quadrotors with suspended payloads or legged robots on uneven terrain, where accurate system modeling is difficult. Our method integrates IMC with large-scale RL training to address these challenges. A simplified predictive model provides a rough yet stable estimate, while large-scale randomized simulations train a robust control policy that compensates for modeling inaccuracies. RL optimizes the policy with the predictive error feedback, revealing external disturbances and nonlinear dynamics, leading to more adaptive and effective control in real-world settings.
@@ III-B Design of a Simplified Predictive Model
The predictive model is key to our framework, enabling discrepancy detection and compensation. Its design faces two challenges: accurately estimating system response and closely aligning predictions with actual behavior. Precise estimation ensures the feedback loop reflects the true robot state, while alignment ensures reliability under varying conditions.
@@
$\tilde{\mathbf{q}}_{k+1}$ $=\mathbf{q}_{k}\otimes\mathbf{q}_{\Delta\theta}$ (5)
$\tilde{\mathbf{\omega}}_{k+1}$ $=\mathbf{\omega}_{k}+\dot{\mathbf{\omega}}_{k}\Delta t$ (6)
We have established an explicit method to predict the next body state, assuming the required motion can be represented by commands at the acceleration or velocity level. Then, given the predicted body state $\tilde{\mathbf{x}}_{k+1}=\left[\tilde{\mathbf{p}}_{k+1},\tilde{\mathbf{q}}_{k+1},\tilde{\mathbf{v}}_{k+1},\tilde{\mathbf{\omega}}_{k+1}\right]$ and the actual body state $\mathbf{x}_{k+1}$, we compute the predictive error $\tilde{\mathbf{d}}_{k+1}$, which is then fed into the policy to provide effective feedback signals.
$\tilde{\mathbf{d}}_{k+1}=\left[\Delta\tilde{\mathbf{p}}_{k+1},1-\cos{\tilde{\Theta}},\Delta\tilde{\mathbf{v}}_{k+1},\Delta\tilde{\mathbf{\omega}}_{k+1}\right]$ (7)
Here, $\Delta\tilde{\mathbf{p}}_{k+1}$, $\Delta\tilde{\mathbf{v}}_{k+1}$, and $\Delta\tilde{\mathbf{\omega}}_{k+1}$ represent the differences in position, velocity, and angular velocity, respectively. We use the cosine distance to quantify the difference between orientations, where $\tilde{\Theta}$ is the angle difference between the two quaternions.
@@ III-C1 Quadrotors
Traditional quadrotor controllers use cascaded structures [38], allowing flexible integration of neural controllers. Recent studies show that using collective thrust and bodyrate (CTBR) as the action space enhances resilience to dynamics mismatches, improves domain transfer, and boosts agility [39, 20, 9]. This design aligns with our framework: the policy’s bodyrate command updates rotation, while collective thrust determines linear acceleration. Consequently, predictive errors, computed from prior policy output and successive body states, integrate directly into the current policy input without modifying the learning algorithm.
@@ III-C2 Quadrupedal robots
Unlike quadrotors, legged robots, particularly quadrupeds and bipeds, are well-suited for RL-based control due to their high-dimensional actuation and terrain interactions [8, 11, 12, 14, 19]. Neural controllers typically generate actuator targets to achieve desired body velocities ($v_{x},v_{y},\omega_{z}$). To integrate our predictive model into the control loop,, we use reference velocity commands as model inputs, deriving linear and angular accelerations by differentiating consecutive commands. Setting other accelerations to zero ensures seamless integration with existing policies. This method is compatible with other algorithms via a plug-and-play feedback loop augmenting policy inputs.
@@ IV Experimental Results
TABLE II: Simulation results for tracking errors (unit: m) with quadrotors, averaged over five random seeds. Bold values denote the best results. Blue cells indicate performance under random external force disturbances, with percentages in brackets representing the relative change in error compared to static conditions. T and S refer to the teacher and student models for RMA, respectively, while DATT with GT indicates the use of ground-truth disturbance inputs instead of $\mathcal{L}_{1}$ estimates.
@@
Train $\mathcal{L}_{1}$-PID 0.214 0.237 (+10.74%) 0.043 0.054 (+25.58%) crashed crashed
$\mathcal{L}_{1}$-MPPI 0.202 0.215 (+6.44%) 0.092 0.117 (+27.17%) 0.193 0.216 (+11.91%) 0.169 0.186 (+10.06%)
CAL 0.087 0.157 (+80.46%) 0.056 0.118 (+110.71%) 0.103 0.175 (+69.90%) 0.105 0.164 (+56.19%)
@@
Eval $\mathcal{L}_{1}$-PID 0.372 0.416 (+11.82%) 0.061 0.107 (+75.41%) crashed crashed
$\mathcal{L}_{1}$-MPPI 0.212 0.234 (+10.38%) 0.097 0.132 (+36.08%) 0.207 0.253 (+22.22%) 0.176 0.218 (+23.86%)
CAL 0.224 0.321 (+43.30%) 0.171 0.251 (+46.78%) 0.262 0.364 (+38.93%) 0.265 0.358 (+35.09%)
@@ IV-B Experiments on Quadrotors
We compare our method for quadrotors against two classical control methods ($\mathcal{L}_{1}$-PID and $\mathcal{L}_{1}$-MPPI) and four learning-based baselines (CAL [28], PPO [27], RMA [12], DATT [20]), using two variants each for RMA and DATT to enable deeper analyses. All methods use the same PPO training parameters. For CAL [28], a sum combination of MPPI and RL, MPPI serves as a sampling-based non-linear MPC method, with the RL policy compensating the MPPI output. We set the sampling number to 4096 and the rollout horizon to 20 in MPPI’s sampling procedure. PPO employs a three-layer MLP policy and a temporal convolutional network (TCN) to process future waypoints in the reference trajectory, which is inherited by all other methods. In addition, RMA adopts a teacher-student training framework, with RMA(T) as the teacher model and RMA(S) as the student. RMA(T) integrates dynamic parameters and ground-truth force disturbances, while RMA(S) uses a TCN to approximate the teacher’s latent state from state-action histories. In contrast, DATT utilizes an $\mathcal{L}_{1}$ disturbance estimator and tracking error inputs to improve trajectory-tracking robustness. To separate the $\mathcal{L}_{1}$ estimator’s effect, we also include a privileged variant, DATT(GT), which directly uses ground-truth force disturbances in place of the estimator.
As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics. Compared to state-of-the-art learning-based methods, classical adaptive control methods perform worse due to their reliance on handcrafted parameters, which fail to generalize across diverse conditions. Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator. However, its performance declines noticeably with unexpected dynamics in evaluation. Comparing DATT(GT) and DATT($\mathcal{L}_{1}$) reveals that the learned RL policy generalizes well when provided with ground-truth inputs, suggesting that the performance drop is primarily due to the handcrafted $\mathcal{L}_{1}$ estimator, which struggles to generalize across different systems. On the other hand. the teacher-student training approach in RMA results in the student policy underperforming compared to the teacher policy and shows instability under unseen systems, highlighting its unreliability for learning-based prediction.
@@ IV-C Experiments on Quadrupedal Robots
For quadrupedal robots, we compare our methods with PPO [27], RMA [12] and HIM [13]. PPO and RMA serve as standard and common baselines in continuous control for quadrupeds, while HIM leverages unsupervised contrastive learning to train a predictive encoder for the next latent state and body velocity. Here, we directly use the teacher-policy of RMA to set a strong baseline. Our method enhances the PPO policy by integrating predictive error feedback. Additionally, we evaluate a variant of our method, adding predictive error feedback to both the inputs of the HIM policy and its predictive encoder during training and evaluation.
As shown in Tab. III, our method outperforms HIM by 26–44% across various test conditions and matches or surpasses the privileged RMA. The combination of our framework with HIM also shows improvements over the original HIM. Interestingly, the predictive encoder in HIM shows a slight degradation compared to standard PPO when combined with our NeuralIMC structure. We hypothesize that this is because the predictive error provides sufficient signal for robust control, rendering the additional impact of predictive encoding negligible. This suggests that predictive error feedback alone is sufficiently informative for enhancing control robustness, simplifying the neural control pipeline, and facilitating integration with other learning algorithms.